# Phonon Induced Contrast in Matter Wave Interferometer

### v1

@@ Appendix A Initial thermal equilibrium mixed state
where $\beta=1/k_{\text{B}}T$. The mathematical calculations from the second line to the third line in Eq. (33) are somewhat complex, but this portion of the content can be found in any material related to thermodynamics and statistical mechanics. The average energy is, thus,
@@
$\left\langle\frac{\dot{u}_{q}^{2}}{2}\right\rangle=\left\langle\frac{\omega_{q}^{2}u_{q}^{2}}{2}\right\rangle=\frac{\langle E\rangle_{q}}{2}=\frac{\hbar\omega}{4}\operatorname{coth}\frac{\hbar\omega}{2T}.$ (35)
@@
Appendix B Displacement and Wigner representation
Here, we supplement some details about Eq. (22). Based on the analysis from Fig. 3, we can consider the coupling between phonon modes after the natural frequency and magnetic field (spin-magnetic field-phonon interaction and phonon-diamagnetic force) as perturbations. Therefore, we can represent the time evolution of the phonon wavefunction in the interaction picture for a specific mode $q$ as,
@@ Appendix B Displacement and Wigner representation
The integral terms in the third line of Eq. (22) precisely represent the differences in phonon amplitudes and momentum given in Eq. (19). Here, one can observe that if the force acting on the atomic chain is coupled with spin and magnetic field, then $V_{\text{I}}^{\text{L}}=-V_{\text{I}}^{\text{R}}$. For diamagnetic interaction, since the diamagnetic force points in the same direction, the situation is different. We use the Wigner representation to calculate the average of (22). under the Wigner representation, the average of an operator $\hat{A}(x)$ can be calculated as,
@@
$\frac{\hbar\omega}{2}\rightarrow\frac{\hbar\omega}{2}\operatorname{coth}\frac{\hbar\omega_{q}}{2k_{\text{B}}T},~{}~{}~{}~{}~{}~{}~{}\frac{\hbar}{2\omega_{q}}\rightarrow\frac{\hbar}{2\omega_{q}}\operatorname{coth}\frac{\hbar\omega_{q}}{k_{\mathrm{B}}T}.$ (43)
Appendix C Contrast loss due to intrinsic dipole-phonon coupling
In this section, we present the contrast loss due to dipole-phonon coupling. It’s worth noting that the dipole-phonon coupling we consider does not involve energy level transitions of the dipole, such as in the Jaynes-Cummings model. We only consider the overlap of the positions and momenta of the phonons on both sides of the interferometer arm, that is, the motion state of the phonons. Here, let’s first consider the case of an intrinsic dipole. In diamond materials embedded with NV centres, due to the NV centre itself carrying a negative charge, the inevitable occurrence of compensatory positive charges arises when it replaces a carbon atom in the material. As a result, a dipole moment $d_{0}$ arises, and its specific value depends on the distance from the compensatory positive charges to the NV centre. This value is approximately one Debye. Here, we aim to study the effect of the dipole on lattice vibrations through its interaction with an external electric field. Therefore, we assume that this intrinsic dipole overlaps with the NV centre, meaning that the compensatory positive charges and the NV centre are located in the same lattice site. Same to the analysis in the main text, let us denote the position of a single intrinsic dipole as $X_{c}(t)+x_{c}$, where $X_{c}(t)$ represents the equilibrium position while $x_{c}$ represent dipole’s vibration. The dipole-phonon interaction Hamiltonian is again mediated by an extra electric field $E(X_{c})$ as.
@@ Appendix C Contrast loss due to intrinsic dipole-phonon coupling
Recall that the diamond has its trajectories as illustrated in Fig. 2, the $X_{c}$ can be expressed explicitly as (13) and $E(X_{c})$ represents the electric field experienced by the dipole. Again, due to the dipole being in a spatial superposition state (considering the diamond entering the SG apparatus), if there is a difference in the derivatives of $E(X_{c})$ on both sides of the interferometer arm, there will be differences in the amplitudes and momentum of the phonons, as illustrated in (19). Obviously, if $E(X_{c})$ is linear, the coupling between the dipole and the electric field will not cause such differences.
Appendix D Contrast loss due to induced dipole-phonon coupling

### v2

@@ Appendix B Initial thermal equilibrium mixed state
where $\beta=1/k_{\text{B}}T$. The calculation from the second to the third line in Eq. (36) can be found in, for example, Kittel and McEuen (2018). The average energy is, thus,
@@
$\left\langle\frac{\dot{u}_{q}^{2}}{2}\right\rangle=\left\langle\frac{\omega_{q}^{2}u_{q}^{2}}{2}\right\rangle=\frac{\langle E\rangle_{q}}{2}=\frac{\hbar\omega}{4}\operatorname{coth}\frac{\hbar\omega}{2k_{\mathrm{B}}T}.$ (38)
@@
Appendix C Overlap and Wigner representation
Here, we supplement some details about Eq. (23). We can represent the time evolution of the phonon wave function in the interaction picture for a specific mode $q$ as,
@@ Appendix C Overlap and Wigner representation
The integral terms in the third line of Eq. (23) precisely represent the differences in phonon amplitudes and momentum given in Eq. (20). Here, one can observe that if the force acting on the atomic chain is coupled with spin and magnetic field, then $V_{\text{I}}^{\text{L}}=-V_{\text{I}}^{\text{R}}$. For diamagnetic interaction, since the diamagnetic force points in the same direction, the situation is different.
We use the Wigner representation to calculate the average of (23). under the Wigner representation, the average of an operator $\hat{A}(x)$ can be calculated as,
@@
$\frac{\hbar\omega}{2}\rightarrow\frac{\hbar\omega}{2}\operatorname{coth}\frac{\hbar\omega_{q}}{2k_{\text{B}}T},~{}~{}~{}~{}~{}~{}~{}\frac{\hbar}{2\omega_{q}}\rightarrow\frac{\hbar}{2\omega_{q}}\operatorname{coth}\frac{\hbar\omega_{q}}{2k_{\mathrm{B}}T}.$ (46)
Appendix D Contrast loss due to intrinsic dipole-phonon coupling
In this section, we present the contrast loss due to dipole-phonon coupling. It’s worth noting that the dipole-phonon coupling we consider does not involve energy level transitions of the dipole, such as in the Jaynes-Cummings model. We only consider the overlap of the positions and momenta of the phonons on both sides of the interferometer arm, that is, the motion state of the phonons.
Here, let’s first consider the case of an intrinsic dipole. In diamond materials embedded with NV centres, due to the NV centre itself carrying a negative charge, the inevitable occurrence of compensatory positive charges arises when it replaces a carbon atom in the material. As a result, a dipole moment $d_{0}$ arises, and its specific value depends on the distance from the compensatory positive charges to the NV centre. This value is approximately one Debye. Here, we aim to study the effect of the dipole on lattice vibrations through its interaction with an external electric field. Therefore, we assume that this intrinsic dipole overlaps with the NV centre, meaning that the compensatory positive charges and the NV centre are located in the same lattice site.
Same to the analysis in the main text, let us denote the position of a single intrinsic dipole as $X_{c}(t)+x_{c}$, where $X_{c}(t)$ represents the equilibrium position while $x_{c}$ represent dipole’s vibration. The dipole-phonon interaction Hamiltonian is again mediated by an extra electric field $E(X_{c})$ as.
@@ Appendix D Contrast loss due to intrinsic dipole-phonon coupling
Recall that the diamond has its trajectories as illustrated in Fig. 2, the $X_{c}$ can be expressed explicitly as (12) and $E(X_{c})$ represents the electric field experienced by the dipole.
Again, due to the dipole being in a spatial superposition state (considering the diamond entering the SG apparatus), if there is a difference in the derivatives of $E(X_{c})$ on both sides of the interferometer arm, there will be differences in the amplitudes and momentum of the phonons, as illustrated in (20). Obviously, if $E(X_{c})$ is linear, the coupling between the dipole and the electric field will not cause such differences.
Appendix E Contrast loss due to induced dipole-phonon coupling